### v1

# MoTCoder: Elevating Large Language Models with Modular of Thought for Challenging Programming Tasks

Abstract
Large Language Models (LLMs) have showcased impressive capabilities in handling straightforward programming tasks. However, their performance tends to falter when confronted with more challenging programming problems. We observe that conventional models often generate solutions as monolithic code blocks, restricting their effectiveness in tackling intricate questions. To overcome this limitation, we present Modular-of-Thought Coder (MoTCoder). We introduce a pioneering framework for MoT instruction tuning, designed to promote the decomposition of tasks into logical sub-tasks and sub-modules. Our investigations reveal that, through the cultivation and utilization of sub-modules, MoTCoder significantly improves both the modularity and correctness of the generated solutions, leading to substantial relative pass@1 improvements of 12.9% on APPS and 9.43% on CodeContests. Our codes are available at https://github.com/dvlab-research/MoTCoder.
[…]
itemsep=0pt,partopsep=0pt,parsep=0pt,topsep=0pt
[…]
1 Introduction
It has been a longstanding objective in the field of artificial intelligence to develop systems capable of generating executable and functionally correct computer programs to address intricate problems (Manna & Waldinger, 1971). Recently, there has been considerable attention on Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2023; Chowdhery et al., 2022; Anil et al., 2023; Hoffmann et al., 2022; Rae et al., 2021; Zeng et al., 2022; Touvron et al., 2023; Zhang et al., 2022), which showcases remarkable success. These models leverage extensive pre-training and undergo further fine-tuning with detailed instruction data (Ouyang et al., 2022), leading to SOTA performance across diverse tasks. Initially designed for natural languages, these models have been expanded through the instruction fine-tuning for code modeling capabilities (Rozière et al., 2023; Black et al., 2021; Chen et al., 2021c). This extension has led to impressive performance in generating code from natural language problem descriptions (Li et al., 2023; Luo et al., 2023a; Wang et al., 2023a; Li et al., 2022b; Nijkamp et al., 2023; Zheng et al., 2023; Fried et al., 2022; Chen et al., 2021a; Wang et al., 2021).
Figure 1: APPS test results by pass@k (%). Comparison of our MoTCoder with previous SOTAs.
However, when confronted with highly intricate coding challenges, the current SOTA models still struggle to match the seasoned developers (Hendrycks et al., 2021b; Li et al., 2022a; Shinn et al., 2023), mainly due to their overly simplistic generation approach. Many previous LLMs produce the code solution as a single monolithic block, instead of breaking down of the task into logical sub-tasks. However, tackling highly complex problems with a single code module is practically unattainable. In contrast, adept developers devise modularized solutions, organizing them into high-level logical sub-modules. Subsequently, they complete and combine these components to efficiently enhance their final solutions.
In our Modular-of-Thought Coder (MoTCoder), our objective is to enhance the modularization capabilities of coding Large Language Models (LLMs) by introducing modular-of-thought prompting. This approach guides LLMs to break down their solutions into modular segments, where each segment represents an abstract function dedicated to a high-level logical sub-task. To train the model to adhere to the MoT prompt, we generate instructional data using a process termed MoT Code Instruction Evolution.
As depicted in Figure 2, our MoT code instruction evolution framework comprises two key stages: In the initial stage, we tailor the evolutionary prompt process. LLMs are instructed to outline necessary sub-modules, generating only their function headers and docstrings that describe their intended usage. Subsequently, the instruction guides the model in implementing these modules and eventually combining them into a comprehensive final solution. In the second stage, we fine-tune the LLM using our created code instruction-following training set, resulting in our MoTCoder model.
Our experiments reveal that MoTCoder significantly enhances LLM performance, establishing new SOTA results on challenging code tasks in APPS (Hendrycks et al., 2021b) and CodeContests  (Li et al., 2022a). Specifically, MoTCoder elevates the pass@1 performance by over 12.9% on APPS and 9.43% on CodeContests.
2 Related Works
Figure 2: The illustration of our MoTCoder training framework. We first perform the modular-of-thought instructions evolution, including the filtration of instructions that do not undergo successful evolution. Subsequently, we delve into the MoT instruction tuning and testing of the model.
2.1 Large Language Models
General LLMs. In recent times, Large Language Models (LLMs) have exhibited remarkable prowess across a wide array of tasks. Leading technology companies have made significant advancements in developing highly proficient closed-source LLMs, including OpenAI’s GPT3 and GPT4 (Brown et al., 2020; OpenAI, 2023), Google’s PaLM (Chowdhery et al., 2022; Anil et al., 2023), Bard [^1] https://bard.google.com/, DeepMind’s Chinchilla (Hoffmann et al., 2022), and Gopher (Rae et al., 2021), as well as Anthropic’s Claude [^2] https://www.anthropic.com/index/introducing-claude. The AI community has also observed the release of several open-source LLMs, where model weights are made publicly available. EleutherAI has contributed GPT-NeoX-20B (Black et al., 2022) and GPT-J-6B (Wang & Komatsuzaki, 2021). Google has released UL2-20B (Tay et al., 2022). Tsinghua University has introduced GLM-130B (Zeng et al., 2022). Meta has released OPT (Zhang et al., 2022) and LLaMA (Touvron et al., 2023).
Coding LLMs. Recent research has introduced a significant number of LLMs tailored for code-related tasks to address the challenges of code understanding and generation. Closed-source models include OpenAI’s Codex (Chen et al., 2021a) and Code-Davinci (Microsoft, 2023). Google has proposed PaLM-Coder (Chowdhery et al., 2022). These models excel on popular code completion benchmarks such as HumanEval (Chen et al., 2021b) and MBPP (Austin et al., 2021). On the open-source front, Salesforce has introduced CodeGen (Nijkamp et al., 2023), CodeT5 (Wang et al., 2021), and CodeT5+ (Wang et al., 2023b). Tsinghua University has contributed CodeGeeX (Zheng et al., 2023), and the BigCode Project has developed StarCoder (Li et al., 2023). These models have demonstrated significant advancements in code-related tasks.
2.2 Instruction Fine-Tuning
General Instruction Tuning. In its early stages, the core aim of instruction fine-tuning was to amplify the cross-task generalization capabilities of Language Models (LMs). This was accomplished by subjecting LMs to fine-tuning using an extensive corpus of public Natural Language Processing (NLP) tasks. Pioneering this approach, T5 (Raffel et al., 2020) underwent training on a diverse set of supervised text-to-text tasks. Subsequent endeavors like FLAN (Wei et al., 2022a), ExT5 (Aribandi et al., 2022), T0 (Sanh et al., 2022), and UnifiedQA (Khashabi et al., 2020) broadened the spectrum of tasks, fortifying the overall generalization capability of LMs. Noteworthy contributions from ZeroPrompt (Xu et al., 2022) and FLAN-T5 (Chung et al., 2022) pushed boundaries by incorporating thousands of tasks into their training pipelines. OpenAI has taken an alternative route by enlisting human annotators to contribute an extensive corpus of human instructions, encompassing diverse formats and a broad spectrum of task types. Building upon this dataset, OpenAI trained its GPT-3 (Brown et al., 2020) model to create InstructGPT (Ouyang et al., 2022), which better aligns with users’ inputs. This developmental trajectory has given rise to notable works such as ChatGPT. In the open-source realm, Alpaca (Taori et al., 2023) adopts the self-instruct method (Wang et al., 2022), leveraging ChatGPT to generate data for training. Vicuna (Chiang et al., 2023) utilizes user-shared conversations collected from ShareGPT.com to train its models. Introducing the Evol-Instruct method, WizardLM (Xu et al., 2023) involves evolving existing instruction data to generate more intricate and diverse datasets. In contrast to these generalized instruction fine-tuning approaches, WizardCoder (Luo et al., 2023b) applies the Evol-Instruct method specifically in the domain of Code LLMs.
Chain-of-Thought Instruction Tuning. While large language models have demonstrated impressive capabilities in handling straightforward programming tasks, their performance tends to diminish when faced with more challenging programming problems. A notable limitation is observed in conventional models, where solutions are often generated as monolithic code blocks, limiting their effectiveness in addressing intricate questions. Drawing inspiration from the success of chain-of-thought (CoT) prompting in enhancing performance in reasoning-based tasks (Zhou et al., 2023; Wei et al., 2022b; Kojima et al., 2022; Le et al., 2023), we introduce a pioneering framework for Modular-of-Thought (MoT) instruction tuning. MoT is designed to facilitate the decomposition of tasks into logical sub-tasks and sub-modules. Our investigations reveal that, through the cultivation and utilization of sub-modules, MoTCoder significantly enhances both the modularity and correctness of the generated solutions.
3 Methods
Figure 3: To evolve conventional instructions into modular-of-thought instructions, our MoT instruction evolution approach guides the models through a two-step process. Initially, the models are instructed to outline the necessary sub-modules, generating only their function headers and docstrings that describe their intended usage. The subsequent instruction then guides the model to implement these modules and integrate them into a cohesive final solution. This instruction is complemented by a one-shot example to further enhance understanding.
The depiction of our MoTCoder training framework is presented in  Figure 2. In the initial stage, we introduce the modular-of-thought instructions evolution and evolution examination in  Section 3.1. Following this, we elaborate on the MoT instruction tuning and testing of the model in  Section 3.2.
3.1 Modular-of-Thought Instruction Evolution
In this section, we propose the modular-of-thought instruct evolution approach. We initiate the evolution process with a given initial instruction dataset $D^{(0)}={(I_{n}^{(0)},\hat{O}_{n}^{(0)})}_{1\leq n\leq N}$, where $I_{n}^{(0)}$ represents the $n$-th instruction in $D^{(0)}$, $O_{n}^{(0)}$ is the corresponding output, and $N$ is the number of samples in $D^{(0)}$. In each evolution step, we upgrade all $I^{(k)}$ in $D^{(k)}$ to $I^{(k+1)}$ by employing an LLM instruction evolution prompt. Subsequently, we use the LLM to generate corresponding outputs $\hat{O}^{(k+1)}$ for the newly evolved $I^{(k+1)}$. Thus, we obtain an evolved instruction dataset $\{D^{(0)}\dots D^{(k+1)}\}$. By iteratively conducting $K$ evolutions, we sequentially acquire the final MoT instruction dataset $\{D^{(0)}\dots D^{(K)}\}$.
Our process for the evolution of modular-of-thought instructions comprises two key stages: MoT instructions evolution, and evolution examination, involving the filtration of instructions that do not successfully evolve. We commence by populating the instruction pool with the provided initial instruction dataset $D^{(0)}$. In each subsequent evolution epoch, normal instructions from the preceding epoch are retrieved from the pool. Subsequently, we employ the instruction evolver to advance each retrieved instruction and the instruction eliminator to evaluate whether the evolution is unsuccessful. Instructions that evolve successfully are integrated into the final pool, while unsuccessful ones are returned as they are, with the anticipation of achieving successful enhancement in the following evolution epoch.
Normal Instruction.
In general, a code sequence generated by a language model $\theta$ through the autoregressive sampling of tokens $\hat{o}_{t}$ is from the parameterized conditional distribution:
$\hat{o}_{t}\sim p_{\theta}(.|\hat{o}_{1:t-1},I),$ (1)
where $I$ represents the input instruction and $\hat{o}_{t}$ is the $t$-th of the flattened output sequence.
Modular-of-Thought Instruction.
Referring to the recent works in code revision (Le et al., 2023), our proposed methodology aims to evolute normal instructions into a sequential code generation process by leading the models through a two-step procedure.
  1. 1.
    Sub-modules. Initially, the models are instructed to outline the required sub-modules, generating only their function headers and docstrings describing their intended usage.
    $\hat{S_{i}}\sim p_{\theta}(.|\hat{S}_{1:i-1},I),$ (2)
    where $\hat{S}_{i}$ represents the $i$-th sub-module outlined by the model and $I$ represents the input instruction.
  2. 2.
    Final solution. The subsequent instruction guides the model to implement these modules and eventually combine them into a comprehensive final solution.
    $\hat{o}_{t}\sim p_{\theta}(.|\hat{o}_{1:t-1},\{\hat{S}_{i}\},I),$ (3)
    where $\hat{o}_{t}$ is the $t$-th of the flattened output sequence.
Modular-of-Thought Instruction Illustration.
The instruction is supplemented with a one-shot example, serving to prompt the model to adhere to the MoT instruction generation strategy. An illustration of the instruction prompt is presented in Figure 3 and further detailed in the Appendix. The instruction generated by MoTCoder encourages the model to decompose a program into sub-modules. This mirrors the methodology commonly employed by developers when addressing intricate coding tasks, where they systematically break down solutions into modular components.
Table 1: Experiments of modular ability. We conducted inference using both normal and modular prompts, as demonstrated in Figure 4 and show APPS test results by pass@1 (%). Comparable approaches include StarCoder (Li et al., 2023) and WizardCoder (Luo et al., 2023b) with the same size of 15B.
Model Size Prompt Introductory Interview Competition All
StarCoder 15B Normal 7.25 6.89 4.08 6.40
Modular 0.59 2.87 1.00 2.04
WizardCoder 15B Normal 26.04 4.21 0.81 7.90
Modular 0.39 0.46 0.66 0.81
Normal 33.80 19.70 11.09 20.80
MoTCoder 15B Modular 35.86 18.79 11.82 20.81
Table 2: Modular-of-thought instruction tuning ablation experiments. We compare the one-step and two-step modular-of-thought instruction tuning strategies. Models are inferred using both normal and modular prompts. APPs test results by pass@1 (%) are shown.
Model Size Prompt Introductory Interview Competition All
Normal 30.76 17.16 11.16 18.68
1-Step MoTCoder 15B Modular 31.69 17.37 11.47 19.05
Normal 33.80 19.70 11.09 20.80
2-Step MoTCoder 15B Modular 35.86 18.79 11.82 20.81
Table 3: Training Datasets. We compare MoTCoder training on only modular dataset and the mixture of modular and normal dataset. APPs test results by pass@1 (%) are shown.
Model Introductory Interview Competition All
Baseline 7.25 6.89 4.08 6.40
Modular Only 31.34 15.16 10.25 17.42
Mixture 33.80 19.70 11.09 20.80
(a) Modular Prompt
(b) Normal Prompt
Figure 4: Example of modular and normal code generation prompts. The modular instructions provide sub-modules including their function headers and docstrings outlining their intended usage. The final resolution executes these modules and then incorporates them into a comprehensive solution.
Evolution Examination.
We categorize the following scenarios as instances of instruction evolution failure:
  1. 1.
    The evolved instruction deviates from the modular-of-thought generation strategy, failing to follow the approach of initially generating sub-modules and then generating the main code.
  2. 2.
    During the first phase, namely the sub-module generation stage, either no sub-modules are generated, or global code is produced.
  3. 3.
    In the second phase, or the main code generation stage, there is either a lack of main code generation or the production of multiple main code segments.
Following the evolution process, we created a modular-of-thought instruction dataset comprising around 24k instructions. The dataset exhibits a distribution ratio of 45:55 between MoT and normal instruction data.
3.2 Modular-of-Thought Training and Inference
We present the MoT prompt training methodology through the introduction of a modular training prompt, as illustrated in Figure 4a. The instructions first specify sub-modules with their function headers and docstrings detailing their intended usage. The ultimate solution involves implementing these modules and subsequently integrating them into a final solution. The {input} in prompts is the problem description and the model is trained to complete the {response} in prompts. We conducted instruction tuning using the base model WizardCoder (Luo et al., 2023b) on our MoT instruction dataset for three epochs. This process took approximately 28 hours using 8 A800 GPUs.
Table 4: APPS test results by pass@k (%). Competitive approaches include the fine-tuned GPT-Neo (Hendrycks et al., 2021a), one-shot Codex (Chen et al., 2021c), fine-tuned AlphaCode (Li et al., 2022c), one-shot StarCoder (Li et al., 2023), one-shot WizardCoder (Luo et al., 2023b) and two-shot CodeLlama series models (Roziere et al., 2023).
Model Size Pass@ Introductory Interview Competition All
GPT-Neo 2.7B 1 3.90 0.57 0.00 1.12
5 5.50 0.80 0.00 1.58
Codex 12B 1 4.14 0.14 0.02 0.92
5 9.65 0.51 0.09 2.25
1000 25.02 3.70 3.23 7.87
AlphaCode 1B 1000 17.67 5.24 7.06 8.09
AlphaCode (Filtered 1k) 5 14.36 5.63 4.58 7.17
AlphaCode (Filtered 10k) 5 18.18 8.21 6.65 9.89
AlphaCode (Filtered 50k) 5 20.36 9.66 7.75 11.42
StarCoder 15B 1 7.25 6.89 4.08 6.40
WizardCoder 15B 1 26.04 4.21 0.81 7.90
CodeLlama 7B 5 10.76 2.01 0.77 3.51
10 15.59 3.12 1.41 5.27
100 33.52 9.40 7.13 13.77
13B 5 23.74 5.63 2.05 8.54
10 30.19 8.12 3.35 11.58
100 48.99 18.40 11.98 23.23
34B 5 32.81 8.75 2.88 12.39
10 38.97 12.16 4.69 16.03
100 56.32 24.31 15.39 28.93
CodeLlama-Python 7B 5 12.72 4.18 1.31 5.31
10 18.50 6.25 2.24 7.90
100 38.26 14.94 9.12 18.44
13B 5 26.33 7.06 2.79 10.06
10 32.77 10.03 4.33 13.44
100 51.60 21.46 14.60 26.12
34B 5 28.94 7.80 3.45 11.16
10 35.91 11.12 5.53 14.96
100 54.92 23.90 16.81 28.69
CodeLlama-Instruct 7B 5 12.85 2.07 1.13 4.04
10 17.86 3.12 1.95 5.83
100 35.37 9.44 8.45 14.43
13B 5 24.01 6.93 2.39 9.44
10 30.27 9.58 3.83 12.57
100 48.73 19.55 13.12 24.10
34B 5 31.56 7.86 3.21 11.67
10 37.80 11.08 5.12 15.23
100 55.72 22.80 16.38 28.10
MoTCoder 15B 1 33.80 19.70 11.09 20.80
code-davinci-002 - 1 29.30 6.40 2.50 10.20
GPT3.5 - 1 48.00 19.42 5.42 22.33
Table 5: APPS test results with code-revision by pass@1 (%). Revision indicates the rounds of revision after the first round of direct generation. Baselines include CodeT5 (Wang et al., 2021), text-davinci and WizardCoder (Luo et al., 2023b) and code-revision methods contain Self-edit (Zhang et al., 2023), CodeRL (Le et al., 2022), Self-repair (Olausson et al., 2023), and CodeChain (Le et al., 2023).
Model Size Revision Introductory Interview Competition All
CodeT5 770M - 6.60 1.03 0.30 2.00
CodeRL+CodeT5 770M 5 7.08 1.86 0.75 2.69
text-davinci-002 - - - - - 7.48
Self-edit+text-davinci-002 - 5 - - - 7.94
WizardCoder 15B - 26.04 4.21 0.81 7.90
CodeChain+WizardCoder 15B 5 26.29 7.49 3.75 10.50
MoTCoder 15B - 33.80 19.70 11.09 20.80
4 Experiments
We showcase the effectiveness of our MoTCoder in addressing challenging code generation tasks, specifically focusing on the following prominent benchmarks:
APPS (Hendrycks et al., 2021b) is a description-to-code generation benchmark from competitive programming platforms Codewars [^3] https://www.codewars.com/, AtCoder [^4] https://atcoder.jp/, Kattis [^5] https://open.kattis.com/, Codeforces [^6] https://codeforces.com/, etc. Building upon prior research (Hendrycks et al., 2021b; Chen et al., 2021c; Li et al., 2022a), we conducted an assessment of the models using the passing rate metric pass@k. This metric is defined as the proportion of problems successfully solved by employing k generated programs for each problem.
CodeContests (Li et al., 2022a) is a competitive programming dataset sourced from Aizu [^7] https://judge.u-aizu.ac.jp, AtCoder [^8] https://atcoder.jp, CodeChef [^9] https://www.codechef.com, Codeforces, HackerEarth [^10] https://www.hackerearth.com, etc. Building upon prior research (Hendrycks et al., 2021b; Chen et al., 2021c; Li et al., 2022a), we conducted an assessment of the models using the passing rate metric pass@k.
4.1 Examination of Modular Ability
In this section, we evaluate the modular capabilities of our model through both normal and modular inference, utilizing prompts as illustrated in Figure 4. For modular prompts, we instruct the model to utilize provided submodules to complete the final code. Conventional models lack this ability to leverage submodules, making the logical and modular aspects more challenging for the model.
Results are in Table 1. On one hand, our MoTCoder demonstrates a significant performance boost compared to models with the same parameters trained with traditional approaches, achieving a remarkable 12.9% improvement on APPs. On the other hand, when employing modular reasoning on models that have not undergone MoT training, there is a considerable decrease in performance. In contrast, our outcomes are minimally affected. This confirms the superior modular capabilities of our approach, which is crucial in large-scale code development environments, where developers often leverage and integrate code sub-modules to construct code frameworks, rather than crafting each line of code from scratch. MoTCoder enables the construction of complete code using only a standardized description of functions, unlocking the potential to generate extensive and complex code frameworks.
The results of MoTCoder also demonstrates that the absence of submodule inputs has a negligible impact on the results. This implies that our method autonomously learns to design submodules without requiring additional hints. In subsequent experiments, we use normal prompts during the inference process for MoTCoder and other methods to ensure a fair comparison.
4.2 Ablation Experiments
In this section, we perform ablation experiments to examine the effectiveness of our MoTCoder training framework.
4.2.1 Modular-of-Thought Instruction-Tuning
We propose two instruction-tuning strategies:
  1. 1.
    One-step strategy. The models directly generate the necessary modules and incorporate them into a complete solution.
  2. 2.
    Two-step strategy. Initially, the instructions outline the essential sub-modules, generating their function headers and docstrings that describe their intended purpose. Subsequently, the instruction guides the model to implement these outlined modules and eventually merge them into a cohesive final solution.
Note that both strategies generate substeps within a single process. As indicated in Table 2, the suggested two-step modular-of-thought instruction tuning strategy outperforms the other both for normal and modular inference. This aligns with the problem-solving approach employed by adept developers: When faced with a problem, skilled developers design solutions with a modular structure, organizing them into high-level logical sub-modules. They subsequently finalize and integrate these components to effectively produce their final solutions.
4.2.2 Training Dataset
In this section, we conduct ablation experiments on the dataset, as illustrated in Table 3. Initially, we finetuned our model on the modular dataset obtained after modular-of-though instruction evolution, resulting in a substantial improvement (11.0%) compared to the baseline of normal fine-tuning. Subsequently, we explore the incorporation of a normal dataset, conducting experiments on a mixture of normal and modular datasets, revealing further performance enhancement of 3.38%. This underscores the significance of data diversity: While we encourage the model to decompose tasks into submodules, we also allow the model to employ direct methods to solve problems. This provides a larger solution space, leading to significant improvements, especially for easy and medium-level problems (2.46% and 4.54%, respectively). The improvement is more modest for difficult problems (0.84%), aligning with the intuition that direct methods are better suited for simpler problems.
Table 6: CodeContests test and validation results by pass@k (%). The rounds of revision after the first round of direct generation is 5 for CodeChain (Le et al., 2023). SOTA models includes code-davinci, GPT3.5 and WizardCoder (Luo et al., 2023b), with code-revision method CodeChain (Le et al., 2023) and filtering method CoT(Chen et al., 2023).
Model Size Revision Validation Test Average
pass@1 pass@5 pass@1 pass@5 pass@1 pass@5
code-davinci-002 - - - - 1.00 - 1.00 -
code-davinci-002 + CodeT - - - - 3.20 - 3.20 -
WizardCoder 15B - 1.11 3.18 1.98 3.27 1.55 3.23
WizardCoder + CodeChain 15B 5 2.35 3.29 2.48 3.30 2.42 3.30
MoTCoder 15B - 2.39 7.69 6.18 12.73 4.29 10.21
GPT3.5 - - 6.81 16.23 5.82 11.16 6.32 13.70
4.3 Results
In this section, we present a comprehensive comparison of our approach with existing large coding model baselines on APPS  (Hendrycks et al., 2021a) in Section 4.3.1 and CodeContest (Li et al., 2022a) in Section 4.3.2. We employ normal inference prompts for all methods.
4.3.1 Results on APPS
We conducted a comparison of our approach with existing LLM baselines on APPs (Hendrycks et al., 2021a). All outcomes are computed using raw predictions without being filtered by the test cases provided in the prompt. We juxtapose our findings with GPT-Neo (Hendrycks et al., 2021a), Codex (Chen et al., 2021c), AlphaCode (Li et al., 2022c), StarCoder (Li et al., 2023), WizardCoder (Luo et al., 2023b) and CodeLlama series models (Roziere et al., 2023).
As depicted in the results from the APPS shown in Table 4, we notice that the modular-of-thought training approach leads to an enhancement in code generation capabilities comparing with previous instruction finetuning models. Our MoTCoder exhibits improved performance across all difficulty levels and demonstrates more substantial gains in the interview and competition-level problems, characterized by more intricate solutions. To provide specifics, the pass@1 performance of MoTCoder-15B surpasses the pass@5 performance of CodeLlama-Python-34B, showcasing significant improvements of 11.9% on interview-level problems and 7.6% on competition-level problems, despite fewer model parameters and inference times. Moreover, our pass@1 on competition-level questions surprisingly outperformed the closed-source model GPT3.5 by an impressive margin of 5.7%. This outcome serves as compelling evidence of the efficacy of our approach in addressing intricate problems.
We also conduct a comparative analysis of our approach against previous LLM baselines with code-revision methods as well. Our included baselines contain Codex (Chen et al., 2021c), CodeT5 (Wang et al., 2021), code-davinci, StarCoder (Li et al., 2023), and WizardCoder (Luo et al., 2023b) and code-revision methods contain Self-edit (Zhang et al., 2023), CodeRL (Wang et al., 2021; Le et al., 2022), Self-repair (Olausson et al., 2023), and CodeChain (Le et al., 2023). The results presented in Table 5 illustrate that MoTCoder exhibits notable performance improvements.
Specifically, MoTCoder achieves a 20.80% pass@1, surpassing the current SOTA, CodeChain + WizardCoder, by 10.3%. CodeChain (Le et al., 2023) is an iterative framework for inference that elicits modularized code generation through a chain of self-revisions. In contrast to CodeChain, our approach introduces a guided modular-of-thought framework during training, making it more intrinsic. This leads to improvements in performance. Notably, this achievement is obtained without the need for any additional revision iterations, thereby avoiding any extra associated inference costs.
4.3.2 Results on CodeContests
We conduct a comprehensive evaluation of our approach on CodeContests, benchmarking it against current state-of-the-art models, including code-davinci, GPT3.5, and WizardCoder (Luo et al., 2023b). The competitive models utilize the code-revision method CodeChain (Le et al., 2023) and the filtering method CodeT (Chen et al., 2023). The results, as depicted in Table 6, reveal notable performance enhancements achieved by MoTCoder. Specifically, MoTCoder attains an impressive 12.73% pass@5 on the test set, surpassing WizardCoder with CodeChain by 9.43%. Remarkably, MoTCoder even outperforms GPT3.5 by 1.57%. Importantly, these achievements are realized without the need for additional revision iterations or filtering by test samples, thereby avoiding any additional associated inference costs or the introduction of extra information.
5 Limitations
Through experimentation, it has been observed that although our MoTCoder elevates the model’s modular and logical capabilities, leading to a commendable performance in addressing competitive-level programming challenges, it concurrently sacrifices some of the model’s conversational abilities. This finding, along with corroborative evidence from prior research (Roziere et al., 2023), underscores the persistent and crucial nature of addressing the forgetting problem. We remain committed to ongoing exploration of this issue in our future work.
6 Conclusion
We propose a novel framework, MoTCoder, for tuning instructions based on a modular of thought. MoTCoder directs LLMs to decompose their solutions into modular segments, culminating in the synthesis of a comprehensive final solution. Our experiments demonstrate that modular-of-thought strategy significantly improves LLM performance, achieving SOTA passing rate on challenging code tasks and showcasing superior modular capabilities. It is crucial in large-scale code development environments, since developers commonly leverage code sub-modules to construct code frameworks over crafting each line of code from scratch. MoTCoder empowers the creation of entire code structures using only standardized descriptions of functions. Adopting this methodology, we unlock the potential to generate extensive and intricate code frameworks.

### v2

@@
 Figure 1: APPS test results by pass@k (%). Comparison of our MoTCoder with previous SOTAs.
+Figure 2: The illustration of our MoTCoder training framework.
+We first perform the modular-of-thought instructions evolution, including the filtration of instructions that do not undergo successful evolution.
+Subsequently, we delve into the MoT instruction tuning and testing of the model.
+Figure 3: To evolve conventional instructions into modular-of-thought instructions, our MoT instruction evolution approach guides the models through a two-step process.
+Initially, the models are instructed to outline the necessary sub-modules, generating only their function headers and docstrings that describe their intended usage.
+The subsequent instruction then guides the model to implement these modules and integrate them into a cohesive final solution.
+This instruction is complemented by a one-shot example to further enhance understanding.
 However, when confronted with highly intricate coding challenges, the current SOTA models still struggle to match the seasoned developers (Hendrycks et al., 2021b; Li et al., 2022a; Shinn et al., 2023), mainly due to their overly simplistic generation approach.
@@
 In the second stage, we fine-tune the LLM using our created code instruction-following training set, resulting in our MoTCoder model.
+Figure 4: Illustration of modular code generation involves the model initially delineating the necessary sub-modules, each comprising a function header and a docstring detailing its intended purpose.
+Following this, the model proceeds to comprehensively implement each module in code and incorporates them as integral components of the finalized solution..
 Our experiments reveal that MoTCoder significantly enhances LLM performance, establishing new SOTA results on challenging code tasks in APPS (Hendrycks et al., 2021b) and CodeContests  (Li et al., 2022a).
@@
 2 Related Works
-Figure 2: The illustration of our MoTCoder training framework.
-We first perform the modular-of-thought instructions evolution, including the filtration of instructions that do not undergo successful evolution.
-Subsequently, we delve into the MoT instruction tuning and testing of the model.
 2.1 Large Language Models
@@
 3 Methods
-Figure 3: To evolve conventional instructions into modular-of-thought instructions, our MoT instruction evolution approach guides the models through a two-step process.
-Initially, the models are instructed to outline the necessary sub-modules, generating only their function headers and docstrings that describe their intended usage.
-The subsequent instruction then guides the model to implement these modules and integrate them into a cohesive final solution.
-This instruction is complemented by a one-shot example to further enhance understanding.
 The depiction of our MoTCoder training framework is presented in  Figure 2.
@@
 The instruction is supplemented with a one-shot example, serving to prompt the model to adhere to the MoT instruction generation strategy.
-An illustration of the instruction prompt is presented in Figure 3 and further detailed in the Appendix.
+An illustration of the instruction prompt is presented in Figure 3 and further detailed in Figure 4.
 The instruction generated by MoTCoder encourages the model to decompose a program into sub-modules.
 This mirrors the methodology commonly employed by developers when addressing intricate coding tasks, where they systematically break down solutions into modular components.
+(a) Modular Prompt
+(b) Normal Prompt
+Figure 5: Example of modular and normal code generation prompts.
+The modular instructions provide sub-modules including their function headers and docstrings outlining their intended usage.
+The final resolution executes these modules and then incorporates them into a comprehensive solution.
+Evolution Examination.
+We categorize the following scenarios as instances of instruction evolution failure:
+  1. 1.
+    The evolved instruction deviates from the modular-of-thought generation strategy, failing to follow the approach of initially generating sub-modules and then generating the main code.
+  2. 2.
+    During the first phase, namely the sub-module generation stage, either no sub-modules are generated, or global code is produced.
+  3. 3.
+    In the second phase, or the main code generation stage, there is either a lack of main code generation or the production of multiple main code segments.
+Following the evolution process, we created a modular-of-thought instruction dataset comprising around 24k instructions.
+The dataset exhibits a distribution ratio of 45:55 between MoT and normal instruction data.
+3.2 Modular-of-Thought Training and Inference
+We present the MoT prompt training methodology through the introduction of a modular training prompt, as illustrated in Figure 5a.
+The instructions first specify sub-modules with their function headers and docstrings detailing their intended usage.
+The ultimate solution involves implementing these modules and subsequently integrating them into a final solution.
+The {input} in prompts is the problem description and the model is trained to complete the {response} in prompts.
+We conducted instruction tuning using the base model WizardCoder (Luo et al., 2023b) on our MoT instruction dataset for three epochs.
+This process took approximately 28 hours using 8 A800 GPUs.
 Table 1: Experiments of modular ability.
@@
 Mixture 33.80 19.70 11.09 20.80
-(a) Modular Prompt
-(b) Normal Prompt
-Figure 4: Example of modular and normal code generation prompts.
-The modular instructions provide sub-modules including their function headers and docstrings outlining their intended usage.
-The final resolution executes these modules and then incorporates them into a comprehensive solution.
-Evolution Examination.
-We categorize the following scenarios as instances of instruction evolution failure:
-  1. 1.
-    The evolved instruction deviates from the modular-of-thought generation strategy, failing to follow the approach of initially generating sub-modules and then generating the main code.
-  2. 2.
-    During the first phase, namely the sub-module generation stage, either no sub-modules are generated, or global code is produced.
-  3. 3.
-    In the second phase, or the main code generation stage, there is either a lack of main code generation or the production of multiple main code segments.
-Following the evolution process, we created a modular-of-thought instruction dataset comprising around 24k instructions.
-The dataset exhibits a distribution ratio of 45:55 between MoT and normal instruction data.
-3.2 Modular-of-Thought Training and Inference
-We present the MoT prompt training methodology through the introduction of a modular training prompt, as illustrated in Figure 4a.
-The instructions first specify sub-modules with their function headers and docstrings detailing their intended usage.
-The ultimate solution involves implementing these modules and subsequently integrating them into a final solution.
-The {input} in prompts is the problem description and the model is trained to complete the {response} in prompts.
-We conducted instruction tuning using the base model WizardCoder (Luo et al., 2023b) on our MoT instruction dataset for three epochs.
-This process took approximately 28 hours using 8 A800 GPUs.
 Table 4: APPS test results by pass@k (%).
-Competitive approaches include the fine-tuned GPT-Neo (Hendrycks et al., 2021a), one-shot Codex (Chen et al., 2021c), fine-tuned AlphaCode (Li et al., 2022c), one-shot StarCoder (Li et al., 2023), one-shot WizardCoder (Luo et al., 2023b) and two-shot CodeLlama series models (Roziere et al., 2023).
+Competitive open-source approaches include CodeT5 (Wang et al., 2021), fine-tuned GPT-Neo (Hendrycks et al., 2021a), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), one-shot StarCoder (Li et al., 2023), and one-shot WizardCoder (Luo et al., 2023b).
+We also present results from closed-source models, namely text-davinci, code-davinci, and GPT3.5.
 Model Size Pass@ Introductory Interview Competition All
-GPT-Neo 2.7B 1 3.90 0.57 0.00 1.12
-5 5.50 0.80 0.00 1.58
-Codex 12B 1 4.14 0.14 0.02 0.92
-5 9.65 0.51 0.09 2.25
-1000 25.02 3.70 3.23 7.87
-AlphaCode 1B 1000 17.67 5.24 7.06 8.09
-AlphaCode (Filtered 1k) 5 14.36 5.63 4.58 7.17
-AlphaCode (Filtered 10k) 5 18.18 8.21 6.65 9.89
-AlphaCode (Filtered 50k) 5 20.36 9.66 7.75 11.42
+CodeT5 770M 1 6.60 1.03 0.30 2.00
+GPT-Neo 2.7B 1 14.68 9.85 6.54 10.15
+5 19.89 13.19 9.90 13.87
+GPT-2 0.1B 1 5.64 6.93 4.37 6.16
+5 13.81 10.97 7.03 10.75
+1.5B 1 7.40 9.11 5.05 7.96
+5 16.86 13.84 9.01 13.48
+GPT-3 175B 1 0.57 0.65 0.21 0.55
 StarCoder 15B 1 7.25 6.89 4.08 6.40
 WizardCoder 15B 1 26.04 4.21 0.81 7.90
-CodeLlama 7B 5 10.76 2.01 0.77 3.51
-10 15.59 3.12 1.41 5.27
-100 33.52 9.40 7.13 13.77
-13B 5 23.74 5.63 2.05 8.54
-10 30.19 8.12 3.35 11.58
-100 48.99 18.40 11.98 23.23
-34B 5 32.81 8.75 2.88 12.39
-10 38.97 12.16 4.69 16.03
-100 56.32 24.31 15.39 28.93
-CodeLlama-Python 7B 5 12.72 4.18 1.31 5.31
-10 18.50 6.25 2.24 7.90
-100 38.26 14.94 9.12 18.44
-13B 5 26.33 7.06 2.79 10.06
-10 32.77 10.03 4.33 13.44
-100 51.60 21.46 14.60 26.12
-34B 5 28.94 7.80 3.45 11.16
-10 35.91 11.12 5.53 14.96
-100 54.92 23.90 16.81 28.69
-CodeLlama-Instruct 7B 5 12.85 2.07 1.13 4.04
-10 17.86 3.12 1.95 5.83
-100 35.37 9.44 8.45 14.43
-13B 5 24.01 6.93 2.39 9.44
-10 30.27 9.58 3.83 12.57
-100 48.73 19.55 13.12 24.10
-34B 5 31.56 7.86 3.21 11.67
-10 37.80 11.08 5.12 15.23
-100 55.72 22.80 16.38 28.10
 MoTCoder 15B 1 33.80 19.70 11.09 20.80
+text-davinci-002 - 1 - - - 7.48
 code-davinci-002 - 1 29.30 6.40 2.50 10.20
@@
 4.3.1 Results on APPS
-We conducted a comparison of our approach with existing LLM baselines on APPs (Hendrycks et al., 2021a).
+We conducted a comparison of our approach with existing large language model baselines on APPs (Hendrycks et al., 2021a).
 All outcomes are computed using raw predictions without being filtered by the test cases provided in the prompt.
-We juxtapose our findings with GPT-Neo (Hendrycks et al., 2021a), Codex (Chen et al., 2021c), AlphaCode (Li et al., 2022c), StarCoder (Li et al., 2023), WizardCoder (Luo et al., 2023b) and CodeLlama series models (Roziere et al., 2023).
-As depicted in the results from the APPS shown in Table 4, we notice that the modular-of-thought training approach leads to an enhancement in code generation capabilities comparing with previous instruction finetuning models.
+Our analysis includes a comparison with open-sourced approaches such as CodeT5 (Wang et al., 2021), fine-tuned GPT-Neo (Hendrycks et al., 2021a), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), one-shot StarCoder (Li et al., 2023), and one-shot WizardCoder (Luo et al., 2023b).
+Additionally, we present results from closed-source models, including text-davinci, code-davinci, and GPT3.5.
+As depicted in the results from the APPS shown in Table 4, we notice that the modular-of-thought training approach leads to an enhancement in code generation capabilities compared with previous instruction finetuning models.
 Our MoTCoder exhibits improved performance across all difficulty levels and demonstrates more substantial gains in the interview and competition-level problems, characterized by more intricate solutions.
-To provide specifics, the pass@1 performance of MoTCoder-15B surpasses the pass@5 performance of CodeLlama-Python-34B, showcasing significant improvements of 11.9% on interview-level problems and 7.6% on competition-level problems, despite fewer model parameters and inference times.
+To provide specifics, the pass@1 performance of MoTCoder-15B surpasses WizardCoder-15b by 15.49% on interview-level problems and 10.28% on competition-level problems.
 Moreover, our pass@1 on competition-level questions surprisingly outperformed the closed-source model GPT3.5 by an impressive margin of 5.7%.

### v3

@@
-MoTCoder: Elevating Large Language Models with Modular of Thought for Challenging Programming Tasks
+MoTCoder: Elevating Large Language Models with Module-of-Thought
 Abstract
@@
 We observe that conventional models often generate solutions as monolithic code blocks, restricting their effectiveness in tackling intricate questions.
-To overcome this limitation, we present Modular-of-Thought Coder (MoTCoder).
-We introduce a pioneering framework for MoT instruction tuning, designed to promote the decomposition of tasks into logical sub-tasks and sub-modules.
-Our investigations reveal that, through the cultivation and utilization of sub-modules, MoTCoder significantly improves both the modularity and correctness of the generated solutions, leading to substantial relative pass@1 improvements of 12.9% on APPS and 9.43% on CodeContests.
+To overcome this limitation, we present Module-of-Thought Coder (MoTCoder).
+We introduce a framework for MoT instruction tuning, designed to promote the decomposition of tasks into logical sub-tasks and sub-modules.
+Our investigations reveal that, through the cultivation and utilization of sub-modules, MoTCoder significantly improves both the modularity and correctness of the generated solutions, leading to substantial relative pass@1 improvements of 60% on APPS and 19% on CodeContests.
 Our codes are available at https://github.com/dvlab-research/MoTCoder.
+1 Introduction
+Developing systems that can generate executable and functionally correct computer programs has long been sought after in artificial intelligence (Manna & Waldinger, 1971).
+Recently, Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2023a; Chowdhery et al., 2022; Anil et al., 2023; Hoffmann et al., 2022; Rae et al., 2021; Zeng et al., 2022; Touvron et al., 2023; Zhang et al., 2022) have showcased remarkable success in many problem domains beyond natural language processing, and is poised as a promising approach to tackle code modeling and generation (as well as other coding related tasks) (Rozière et al., 2023; Black et al., 2021; Chen et al., 2021c).
+Through instruction fine-tuning (Li et al., 2023b; Luo et al., 2023b; Wang et al., 2023a; Li et al., 2022b; Nijkamp et al., 2023; Zheng et al., 2023; Fried et al., 2022; Chen et al., 2021a; Wang et al., 2021), LLMs have achieved impressive performance on code generation benchmarks like HumanEval (Chen et al., 2021b) and MBPP (Austin et al., 2021).
+Yet, when confronted with intricate coding problems such as APPS (Hendrycks et al., 2021b) and CodeContests  (Li et al., 2022a), SOTA models struggle to match seasoned developers (Hendrycks et al., 2021b; Li et al., 2022a; Shinn et al., 2023).
+The main culprit is their overly simplistic generation approach.
+Current models produce the code solution as a single monolithic block in an attempt to solve the problem in one shot.
+While feasible for simple tasks, this approach becomes increasingly inefficient to tackle a complex task which naturally entails solving multiple sub-tasks.
+In contrast, adept developers often devise modularized solutions by breaking down the original problem into more approachable components that can be individually and more efficiently solved.
+Following this intuition, recent works (Jiang et al., 2023; Le et al., 2023) propose iterative code inference for code generation.
+They, however, come with added inference costs and offer only marginal performance gains.
+In this work, we propose to more efficiently improve the modularization capability of coding LLMs using Module-of-Thought (MoT) instruction fine-tuning.
+Our approach guides LLMs to break down their solution into modular components, each representing an abstract function dedicated to a logical sub-task.
+To train the model to adhere to the MoT prompt, we generate instructional data using a process termed MoT Code Instruction Transformation.
+In this process, LLMs are instructed to outline necessary modules, generating only their function headers and docstrings that describe their intended usage.
+Subsequently, the instruction guides the model to implement these modules and eventually combine them into the final solution.
+After that, we fine-tune the LLM with our MoT instruction fune-tuning, resulting in our MoTCoder model.
+In summary, our contributions are threefold:
+  1. 1.
+    We propose a 2-step Module-of-Thought Code Instruction Transformation approach that instructs LLMs to generate modularized code solutions, where each module is to address a logical sub-task and accompanied by standardized annotations.
+  2. 2.
+    We develop Module-of-Thought Coder (MoTCoder), a new model that enhances the modularization capabilities of LLMs with MoT Instruction Tuning.
+It breaks down complex problems into manageable, logically coherent sub-modules, improving both comprehensibility and maintainability of the generated code.
+  3. 3.
+    Our experiments demonstrate that MoTCoder establishes new SOTA results on challenging code generation benchmarks such as APPS and CodeContests.
+Specifically, MoTCoder improves the pass@1 performance by significant margins, exemplified by improvements over existing metrics by 60% on APPS and 19% on CodeContests, as illustrated in Section 1.
 […]
-itemsep=0pt,partopsep=0pt,parsep=0pt,topsep=0pt
-[…]
-1 Introduction
-It has been a longstanding objective in the field of artificial intelligence to develop systems capable of generating executable and functionally correct computer programs to address intricate problems (Manna & Waldinger, 1971).
-Recently, there has been considerable attention on Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2023; Chowdhery et al., 2022; Anil et al., 2023; Hoffmann et al., 2022; Rae et al., 2021; Zeng et al., 2022; Touvron et al., 2023; Zhang et al., 2022), which showcases remarkable success.
-These models leverage extensive pre-training and undergo further fine-tuning with detailed instruction data (Ouyang et al., 2022), leading to SOTA performance across diverse tasks.
-Initially designed for natural languages, these models have been expanded through the instruction fine-tuning for code modeling capabilities (Rozière et al., 2023; Black et al., 2021; Chen et al., 2021c).
-This extension has led to impressive performance in generating code from natural language problem descriptions (Li et al., 2023; Luo et al., 2023a; Wang et al., 2023a; Li et al., 2022b; Nijkamp et al., 2023; Zheng et al., 2023; Fried et al., 2022; Chen et al., 2021a; Wang et al., 2021).
-Figure 1: APPS test results by pass@k (%). Comparison of our MoTCoder with previous SOTAs.
-Figure 2: The illustration of our MoTCoder training framework.
-We first perform the modular-of-thought instructions evolution, including the filtration of instructions that do not undergo successful evolution.
-Subsequently, we delve into the MoT instruction tuning and testing of the model.
-Figure 3: To evolve conventional instructions into modular-of-thought instructions, our MoT instruction evolution approach guides the models through a two-step process.
-Initially, the models are instructed to outline the necessary sub-modules, generating only their function headers and docstrings that describe their intended usage.
-The subsequent instruction then guides the model to implement these modules and integrate them into a cohesive final solution.
-This instruction is complemented by a one-shot example to further enhance understanding.
-However, when confronted with highly intricate coding challenges, the current SOTA models still struggle to match the seasoned developers (Hendrycks et al., 2021b; Li et al., 2022a; Shinn et al., 2023), mainly due to their overly simplistic generation approach.
-Many previous LLMs produce the code solution as a single monolithic block, instead of breaking down of the task into logical sub-tasks.
-However, tackling highly complex problems with a single code module is practically unattainable.
-In contrast, adept developers devise modularized solutions, organizing them into high-level logical sub-modules.
-Subsequently, they complete and combine these components to efficiently enhance their final solutions.
-In our Modular-of-Thought Coder (MoTCoder), our objective is to enhance the modularization capabilities of coding Large Language Models (LLMs) by introducing modular-of-thought prompting.
-This approach guides LLMs to break down their solutions into modular segments, where each segment represents an abstract function dedicated to a high-level logical sub-task.
-To train the model to adhere to the MoT prompt, we generate instructional data using a process termed MoT Code Instruction Evolution.
-As depicted in Figure 2, our MoT code instruction evolution framework comprises two key stages: In the initial stage, we tailor the evolutionary prompt process.
-LLMs are instructed to outline necessary sub-modules, generating only their function headers and docstrings that describe their intended usage.
-Subsequently, the instruction guides the model in implementing these modules and eventually combining them into a comprehensive final solution.
-In the second stage, we fine-tune the LLM using our created code instruction-following training set, resulting in our MoTCoder model.
-Figure 4: Illustration of modular code generation involves the model initially delineating the necessary sub-modules, each comprising a function header and a docstring detailing its intended purpose.
-Following this, the model proceeds to comprehensively implement each module in code and incorporates them as integral components of the finalized solution..
-Our experiments reveal that MoTCoder significantly enhances LLM performance, establishing new SOTA results on challenging code tasks in APPS (Hendrycks et al., 2021b) and CodeContests  (Li et al., 2022a).
-Specifically, MoTCoder elevates the pass@1 performance by over 12.9% on APPS and 9.43% on CodeContests.
+figurePass@1 results on CodeContests (left) and APPS (right). Comparison of our MoTCoder with previous SOTAs.
 2 Related Works
@@
 In recent times, Large Language Models (LLMs) have exhibited remarkable prowess across a wide array of tasks.
-Leading technology companies have made significant advancements in developing highly proficient closed-source LLMs, including OpenAI’s GPT3 and GPT4 (Brown et al., 2020; OpenAI, 2023), Google’s PaLM (Chowdhery et al., 2022; Anil et al., 2023), Bard [^1] https://bard.google.com/, DeepMind’s Chinchilla (Hoffmann et al., 2022), and Gopher (Rae et al., 2021), as well as Anthropic’s Claude [^2] https://www.anthropic.com/index/introducing-claude.
+Leading technology companies have made significant advancements in developing highly proficient closed-source LLMs, including OpenAI’s GPT3 (Brown et al., 2020) and GPT4 (OpenAI, 2023a), Google’s PaLM (Chowdhery et al., 2022; Anil et al., 2023), Bard [^1] https://bard.google.com/, DeepMind’s Chinchilla (Hoffmann et al., 2022), and Gopher (Rae et al., 2021), as well as Anthropic’s Claude [^2] https://www.anthropic.com/index/introducing-claude.
 The AI community has also observed the release of several open-source LLMs, where model weights are made publicly available.
@@
 On the open-source front, Salesforce has introduced CodeGen (Nijkamp et al., 2023), CodeT5 (Wang et al., 2021), and CodeT5+ (Wang et al., 2023b).
-Tsinghua University has contributed CodeGeeX (Zheng et al., 2023), and the BigCode Project has developed StarCoder (Li et al., 2023).
+Tsinghua University has contributed CodeGeeX (Zheng et al., 2023), and the BigCode Project has developed StarCoder (Li et al., 2023b).
 These models have demonstrated significant advancements in code-related tasks.
@@
 Pioneering this approach, T5 (Raffel et al., 2020) underwent training on a diverse set of supervised text-to-text tasks.
-Subsequent endeavors like FLAN (Wei et al., 2022a), ExT5 (Aribandi et al., 2022), T0 (Sanh et al., 2022), and UnifiedQA (Khashabi et al., 2020) broadened the spectrum of tasks, fortifying the overall generalization capability of LMs.
+Subsequent endeavors like FLAN (Wei et al., 2022a), ExT5 (Aribandi et al., 2022), T0 (Sanh et al., 2022a), and UnifiedQA (Khashabi et al., 2020) broadened the spectrum of tasks, fortifying the overall generalization capability of LMs.
 Noteworthy contributions from ZeroPrompt (Xu et al., 2022) and FLAN-T5 (Chung et al., 2022) pushed boundaries by incorporating thousands of tasks into their training pipelines.
@@
 Introducing the Evol-Instruct method, WizardLM (Xu et al., 2023) involves evolving existing instruction data to generate more intricate and diverse datasets.
-In contrast to these generalized instruction fine-tuning approaches, WizardCoder (Luo et al., 2023b) applies the Evol-Instruct method specifically in the domain of Code LLMs.
 Chain-of-Thought Instruction Tuning.
-While large language models have demonstrated impressive capabilities in handling straightforward programming tasks, their performance tends to diminish when faced with more challenging programming problems.
-A notable limitation is observed in conventional models, where solutions are often generated as monolithic code blocks, limiting their effectiveness in addressing intricate questions.
-Drawing inspiration from the success of chain-of-thought (CoT) prompting in enhancing performance in reasoning-based tasks (Zhou et al., 2023; Wei et al., 2022b; Kojima et al., 2022; Le et al., 2023), we introduce a pioneering framework for Modular-of-Thought (MoT) instruction tuning.
-MoT is designed to facilitate the decomposition of tasks into logical sub-tasks and sub-modules.
+In contrast to these generalized approaches, Recent works (Luo et al., 2023c; Yue et al., 2023; Chen et al., 2022b; Gunasekar et al., 2023; Haluptzok et al., 2023) apply instruction tuning method in common-sense reasoning (West et al., 2022), text-summarization (Sclar et al., 2022), mathematical reasoning  (Luo et al., 2023a; Yue et al., 2023), tool use (Patil et al., 2023), coding (Luo et al., 2023b), and general-purpose reasoning (Li et al., 2023a; Zelikman et al., 2022).
+Among them,  Yue et al.
+(2023) curates a diverse corpus of mathematics problems with chain-of-thought or program-of-thought (Chen et al., 2022b) annotations for mathematical reasoning analogous to our plans.
+Gunasekar et al.
+(2023) proposed pre-training models on programming textbooks generated synthetically from GPT3.5.
+Haluptzok et al.
+(2023) similarly generates programming puzzles and corresponding solutions from language models.
+2.3 Prompting Techniques
+A prompt technique known as Chain of Thought (CoT) (Wei et al., 2022b) generated intermediate reasoning steps before the answer to address language reasoning tasks.
+Inspired by CoT, a series of prompting works has been proposed.
+Least-to-most prompting (Zhou et al., 2022) reduces a complex problem into a series of sub-problems and then solves the sub-problems in order, adding the answer to the previous sub-problems to the prompt each time solving begins.
+PAL (Gao et al., 2022) and PoT (Chen et al., 2022a) is proposed to generating code as the intermediate reasoning steps, delegating solving to the compiler, thus improving solution accuracy.
+Similar methods are proposed for simple mathematical (Lewkowycz et al., 2022; Wu et al., 2022), commonsense (Sanh et al., 2022b; Madaan et al., 2022), symbolic reasoning (Yao et al., 2023) and code generation problems (Jiang et al., 2023; Le et al., 2023).
+However, these works can only plan code during generation.
+In comparison, our approach introduces a guided module-of-thought framework during training, making it more intrinsic.
 Our investigations reveal that, through the cultivation and utilization of sub-modules, MoTCoder significantly enhances both the modularity and correctness of the generated solutions.
+Figure 1: Left: To transform conventional instructions into module-of-thought instructions, our MoT instruction transformation approach guides the models through a two-step process.
+Initially, the models are instructed to outline the necessary sub-modules, generating only their function headers and docstrings that describe their intended usage.
+The subsequent instruction then guides the model to implement these sub-modules and integrate them into a cohesive final solution.
+This instruction is complemented by a one-shot example to further enhance understanding.
+Right: Examples of instruction, modularized code and solution.
 3 Methods
-The depiction of our MoTCoder training framework is presented in  Figure 2.
-In the initial stage, we introduce the modular-of-thought instructions evolution and evolution examination in  Section 3.1.
-Following this, we elaborate on the MoT instruction tuning and testing of the model in  Section 3.2.
-3.1 Modular-of-Thought Instruction Evolution
-In this section, we propose the modular-of-thought instruct evolution approach.
-We initiate the evolution process with a given initial instruction dataset $D^{(0)}={(I_{n}^{(0)},\hat{O}_{n}^{(0)})}_{1\leq n\leq N}$, where $I_{n}^{(0)}$ represents the $n$-th instruction in $D^{(0)}$, $O_{n}^{(0)}$ is the corresponding output, and $N$ is the number of samples in $D^{(0)}$.
-In each evolution step, we upgrade all $I^{(k)}$ in $D^{(k)}$ to $I^{(k+1)}$ by employing an LLM instruction evolution prompt.
-Subsequently, we use the LLM to generate corresponding outputs $\hat{O}^{(k+1)}$ for the newly evolved $I^{(k+1)}$.
-Thus, we obtain an evolved instruction dataset $\{D^{(0)}\dots D^{(k+1)}\}$.
-By iteratively conducting $K$ evolutions, we sequentially acquire the final MoT instruction dataset $\{D^{(0)}\dots D^{(K)}\}$.
-Our process for the evolution of modular-of-thought instructions comprises two key stages: MoT instructions evolution, and evolution examination, involving the filtration of instructions that do not successfully evolve.
-We commence by populating the instruction pool with the provided initial instruction dataset $D^{(0)}$.
-In each subsequent evolution epoch, normal instructions from the preceding epoch are retrieved from the pool.
-Subsequently, we employ the instruction evolver to advance each retrieved instruction and the instruction eliminator to evaluate whether the evolution is unsuccessful.
-Instructions that evolve successfully are integrated into the final pool, while unsuccessful ones are returned as they are, with the anticipation of achieving successful enhancement in the following evolution epoch.
-Normal Instruction.
+In this section, we first detail the module-of-thought instruction transformation in Section 3.1 and then introduce the module-of-thought instruction tuning strategy for our MoTCoder in Section 3.2.
+3.1 Module-of-Thought Instruction Transformation
+In this section, we first introduce the normal instruction, followed by our module-of-thought instruction and its assessment.
+Figure 2: Illustration of our 2-step Module-of-Thought Instruction Transformation: Plain code utilizes a single module to directly generate code.
+In contrast, our Module-of-Thought Instruction Transformation first instructs LLMs to outline necessary sub-modules, generating only their function headers and docstrings that describe their intended usage.
+Subsequently, the instructions guide the model in implementing these sub-modules and eventually combining them into a comprehensive final solution.
+3.1.1 Normal Instruction
 In general, a code sequence generated by a language model $\theta$ through the autoregressive sampling of tokens $\hat{o}_{t}$ is from the parameterized conditional distribution:
-$\hat{o}_{t}\sim p_{\theta}(.|\hat{o}_{1:t-1},I),$ (1)
-where $I$ represents the input instruction and $\hat{o}_{t}$ is the $t$-th of the flattened output sequence.
-Modular-of-Thought Instruction.
-Referring to the recent works in code revision (Le et al., 2023), our proposed methodology aims to evolute normal instructions into a sequential code generation process by leading the models through a two-step procedure.
+$o_{t}\sim p_{\theta}(.|o_{1:t-1},I),$ (1)
+where $I$ represents the input instruction and $o_{t}$ is the $t$-th token of the flattened output sequence.
+3.1.2 Module-of-Thought Instruction
+Considering programming encompasses both detailed coding abilities and broader logical or strategic planning capabilities, we propose a transformation pipeline that enhances code clarity and structure targeting detailed coding proficiencies and integrating descriptions based on natural language for the broader strategic planning capabilities.
+According to previous researches (Jain et al., 2023), straightforward and unambiguous, detailed instructions enhance the model’s efficacy and precision in executing the desired tasks.
+Hence, for intricate data transformation tasks, decomposing the task into simpler, sequential steps yields better outcomes.
+Therefore, our proposed methodology aims to transform normal instructions into a sequential code generation process by leading the models through a two-step procedure, as illustrated in Figure 2 .
   1. 1.
@@
   2. 2.
-    Final solution. The subsequent instruction guides the model to implement these modules and eventually combine them into a comprehensive final solution.
+    Final solution. The subsequent instruction guides the model to implement these sub-modules and eventually combine them into a comprehensive final solution.
     $\hat{o}_{t}\sim p_{\theta}(.|\hat{o}_{1:t-1},\{\hat{S}_{i}\},I),$ (3)
     where $\hat{o}_{t}$ is the $t$-th of the flattened output sequence.
-Modular-of-Thought Instruction Illustration.
 The instruction is supplemented with a one-shot example, serving to prompt the model to adhere to the MoT instruction generation strategy.
-An illustration of the instruction prompt is presented in Figure 3 and further detailed in Figure 4.
-The instruction generated by MoTCoder encourages the model to decompose a program into sub-modules.
+An illustration of the instruction prompt is presented in Figure 1.
+The instruction encourages the model to decompose a program into sub-modules.
 This mirrors the methodology commonly employed by developers when addressing intricate coding tasks, where they systematically break down solutions into modular components.
-(a) Modular Prompt
-(b) Normal Prompt
-Figure 5: Example of modular and normal code generation prompts.
-The modular instructions provide sub-modules including their function headers and docstrings outlining their intended usage.
-The final resolution executes these modules and then incorporates them into a comprehensive solution.
-Evolution Examination.
-We categorize the following scenarios as instances of instruction evolution failure:
+3.1.3 Instruction Assessment
+Throughout these transformations, we introduce guidelines for reviewing our transformed module-of-thought code. We identify the situations below as markers of instruction refinement failure:
   1. 1.
-    The evolved instruction deviates from the modular-of-thought generation strategy, failing to follow the approach of initially generating sub-modules and then generating the main code.
+    The refined instruction diverges from the module-of-thought generation strategy, not adhering to the protocol of initial sub-module creation followed by the main code development.
   2. 2.
-    During the first phase, namely the sub-module generation stage, either no sub-modules are generated, or global code is produced.
+    At the sub-module creation phase, if no sub-modules are formed or if overarching code is developed instead.
   3. 3.
-    In the second phase, or the main code generation stage, there is either a lack of main code generation or the production of multiple main code segments.
-Following the evolution process, we created a modular-of-thought instruction dataset comprising around 24k instructions.
-The dataset exhibits a distribution ratio of 45:55 between MoT and normal instruction data.
-3.2 Modular-of-Thought Training and Inference
-We present the MoT prompt training methodology through the introduction of a modular training prompt, as illustrated in Figure 5a.
-The instructions first specify sub-modules with their function headers and docstrings detailing their intended usage.
-The ultimate solution involves implementing these modules and subsequently integrating them into a final solution.
-The {input} in prompts is the problem description and the model is trained to complete the {response} in prompts.
-We conducted instruction tuning using the base model WizardCoder (Luo et al., 2023b) on our MoT instruction dataset for three epochs.
+    During the main code development phase, the absence of main code creation or the emergence of multiple main code blocks indicates a problem.
+  4. 4.
+    The presence of test cases in the dataset that the transformed program fails to pass. This criterion ensures the transformed programs preserve functional equivalence with the original codes.
+3.2 Module-of-Thought Instruction Tuning
+In this section, we first introduce our module-of-thought dataset, followed by our MoT instruction tuning details.
+3.2.1 MoT Dataset
+We use data sourced from a small part of the training set of APPS (Hendrycks et al., 2021b) and CodeContests (Li et al., 2022a) as our initial set.
+For each problem, we take at most one answer, whereas the original dataset may contain dozens of answers.
+We then use GPT3.5 to generate transformed instructions, with the instruction transformation prompt as shown in Figure 1.
+The generated module-of-thought instruction dataset comprises around 24k instructions.
+We merge the docstrings that describe the intended usage and input-output variables type of the functions with their implementation.
+Some examples from the MoT dataset are in Appendix B.
+3.2.2 MoT Instruction Tuning
+Our MoTCoder underwent instruction tuning utilizing the SOTA coding model Deepseek-Cdoer-Base (Balog et al., 2016) across three epochs on our proposed MoT instruction dataset.
+The model’s maximum input length is 2048 tokens.
+We used a training batch size of 16 and an evaluation batch size of 1 per device, with gradient accumulation steps set to 4.
+The learning rate was initialized at $2\times 10^{-5}$, with a warmup period of 10 steps to gradually adapt the learning rate, following a cosine learning rate scheduler for optimization.
+Additionally, we enabled gradient checkpointing, BF16 precision and more deepseed settings to enhance training efficiency and reduce memory consumption: Our model was configured with the AdamW optimizer and utilized the WarmupLR scheduler to manage the learning rate adjustments effectively.
+To optimize memory and compute resources, we employed a third-stage zero optimization setting and enable communication overlap, contiguous gradients, and large sub group size of $1\times 10^{9}$ to streamline the training process.
+In addition, the maximum live parameters, maximum reuse distance, and the parameter settings for gathering 16-bit weights during model saving were all set to $1\times 10^{9}$.
 This process took approximately 28 hours using 8 A800 GPUs.
-Table 1: Experiments of modular ability.
-We conducted inference using both normal and modular prompts, as demonstrated in Figure 5 and show APPS test results by pass@1 (%).
-Comparable approaches include StarCoder (Li et al., 2023) and WizardCoder (Luo et al., 2023b) with the same size of 15B.
-Model Size Prompt Introductory Interview Competition All
-StarCoder 15B Normal 7.25 6.89 4.08 6.40
-Modular 0.59 2.87 1.00 2.04
-WizardCoder 15B Normal 26.04 4.21 0.81 7.90
-Modular 0.39 0.46 0.66 0.81
-Normal 33.80 19.70 11.09 20.80
-MoTCoder 15B Modular 35.86 18.79 11.82 20.81
-Table 2: Modular-of-thought instruction tuning ablation experiments.
-We compare the one-step and two-step modular-of-thought instruction tuning strategies.
-Models are inferred using both normal and modular prompts.
-APPs test results by pass@1 (%) are shown.
-Model Size Prompt Introductory Interview Competition All
-Normal 30.76 17.16 11.16 18.68
-1-Step MoTCoder 15B Modular 31.69 17.37 11.47 19.05
-Normal 33.80 19.70 11.09 20.80
-2-Step MoTCoder 15B Modular 35.86 18.79 11.82 20.81
-Table 3: Training Datasets. We compare MoTCoder training on only modular dataset and the mixture of modular and normal dataset. APPs test results by pass@1 (%) are shown.
-Model Introductory Interview Competition All
-Baseline 7.25 6.89 4.08 6.40
-Modular Only 31.34 15.16 10.25 17.42
-Mixture 33.80 19.70 11.09 20.80
-Table 4: APPS test results by pass@k (%).
-Competitive open-source approaches include CodeT5 (Wang et al., 2021), fine-tuned GPT-Neo (Hendrycks et al., 2021a), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), one-shot StarCoder (Li et al., 2023), and one-shot WizardCoder (Luo et al., 2023b).
-We also present results from closed-source models, namely text-davinci, code-davinci, and GPT3.5.
-Model Size Pass@ Introductory Interview Competition All
-CodeT5 770M 1 6.60 1.03 0.30 2.00
-GPT-Neo 2.7B 1 14.68 9.85 6.54 10.15
-5 19.89 13.19 9.90 13.87
-GPT-2 0.1B 1 5.64 6.93 4.37 6.16
-5 13.81 10.97 7.03 10.75
-1.5B 1 7.40 9.11 5.05 7.96
-5 16.86 13.84 9.01 13.48
-GPT-3 175B 1 0.57 0.65 0.21 0.55
-StarCoder 15B 1 7.25 6.89 4.08 6.40
-WizardCoder 15B 1 26.04 4.21 0.81 7.90
-MoTCoder 15B 1 33.80 19.70 11.09 20.80
-text-davinci-002 - 1 - - - 7.48
-code-davinci-002 - 1 29.30 6.40 2.50 10.20
-GPT3.5 - 1 48.00 19.42 5.42 22.33
-Table 5: APPS test results with code-revision by pass@1 (%).
-Revision indicates the rounds of revision after the first round of direct generation.
-Baselines include CodeT5 (Wang et al., 2021), text-davinci and WizardCoder (Luo et al., 2023b) and code-revision methods contain Self-edit (Zhang et al., 2023), CodeRL (Le et al., 2022), Self-repair (Olausson et al., 2023), and CodeChain (Le et al., 2023).
-Model Size Revision Introductory Interview Competition All
-CodeT5 770M - 6.60 1.03 0.30 2.00
-CodeRL+CodeT5 770M 5 7.08 1.86 0.75 2.69
-text-davinci-002 - - - - - 7.48
-Self-edit+text-davinci-002 - 5 - - - 7.94
-WizardCoder 15B - 26.04 4.21 0.81 7.90
-CodeChain+WizardCoder 15B 5 26.29 7.49 3.75 10.50
-MoTCoder 15B - 33.80 19.70 11.09 20.80
 4 Experiments
-We showcase the effectiveness of our MoTCoder in addressing challenging code generation tasks, specifically focusing on the following prominent benchmarks:
+We demonstrate the efficacy of MoTCoder in tackling intricate code-generation tasks, rather than those that can be solved with just a few lines, as exemplified by HumanEval (Chen et al., 2021b) and MBPP (Austin et al., 2021).
+Specifically, we focus on the following prominent benchmarks.
 APPS (Hendrycks et al., 2021b) is a description-to-code generation benchmark from competitive programming platforms Codewars [^3] https://www.codewars.com/, AtCoder [^4] https://atcoder.jp/, Kattis [^5] https://open.kattis.com/, Codeforces [^6] https://codeforces.com/, etc.
@@
 Building upon prior research (Hendrycks et al., 2021b; Chen et al., 2021c; Li et al., 2022a), we conducted an assessment of the models using the passing rate metric pass@k.
-4.1 Examination of Modular Ability
-In this section, we evaluate the modular capabilities of our model through both normal and modular inference, utilizing prompts as illustrated in Figure 5.
-For modular prompts, we instruct the model to utilize provided submodules to complete the final code.
-Conventional models lack this ability to leverage submodules, making the logical and modular aspects more challenging for the model.
-Results are in Table 1.
-On one hand, our MoTCoder demonstrates a significant performance boost compared to models with the same parameters trained with traditional approaches, achieving a remarkable 12.9% improvement on APPs.
-On the other hand, when employing modular reasoning on models that have not undergone MoT training, there is a considerable decrease in performance.
-In contrast, our outcomes are minimally affected.
-This confirms the superior modular capabilities of our approach, which is crucial in large-scale code development environments, where developers often leverage and integrate code sub-modules to construct code frameworks, rather than crafting each line of code from scratch.
-MoTCoder enables the construction of complete code using only a standardized description of functions, unlocking the potential to generate extensive and complex code frameworks.
-The results of MoTCoder also demonstrates that the absence of submodule inputs has a negligible impact on the results.
-This implies that our method autonomously learns to design submodules without requiring additional hints.
-In subsequent experiments, we use normal prompts during the inference process for MoTCoder and other methods to ensure a fair comparison.
+Table 1: Training Datasets. We compare MoTCoder training on normal and modular-of-thought dataset. APPs test results by pass@1 (%) are shown.
+Model Introductory Interview Competition All
+Normal 50.20 26.65 11.52 28.33
+Module-of-Thought 50.01 29.81 14.36 30.76
+Table 2: APPS test results by pass@k (%).
+Competitive approaches include CodeT5 (Wang et al., 2021), fine-tuned GPT-Neo (Hendrycks et al., 2021a), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), one-shot StarCoder (Li et al., 2023b), WizardCoder (Luo et al., 2023c), CodeLlama series (Roziere et al., 2023), text-davinci, GPT-4 (OpenAI, 2023b).
+We also include code-revision methods containing Self-edit (Zhang et al., 2023), CodeRL (Le et al., 2022), Self-repair (Olausson et al., 2023), and CodeChain (Le et al., 2023).
+Model Size Introductory Interview Competition All
+CodeT5 770M 6.60 1.03 0.30 2.00
+CodeRL+CodeT5 770M 7.08 1.86 0.75 2.69
+text-davinci-002 - - - - 7.48
+Self-edit+text-davinci-002 - - - - 7.94
+GPT-2 0.1B 5.64 6.93 4.37 6.16
+1.5B 7.40 9.11 5.05 7.96
+GPT-Neo 2.7B 14.68 9.85 6.54 10.15
+GPT-3 175B 0.57 0.65 0.21 0.55
+StarCoder 15B 7.25 6.89 4.08 6.40
+WizardCoder 15B 26.04 4.21 0.81 7.90
+CodeChain+WizardCoder 15B 26.29 7.49 3.75 10.50
+Octocoder 16B 16.50 7.92 4.61 8.97
+Codellama 7B 14.15 6.63 4.00 7.61
+13B 23.94 13.50 9.80 14.85
+34B 32.01 18.61 10.19 19.61
+Codellama-Python 7B 18.83 8.62 4.47 9.83
+13B 26.40 13.44 6.86 14.72
+34B 26.45 16.61 8.77 17.01
+Codellama-Instruct 7B 14.20 6.63 4.43 7.70
+13B 22.41 14.34 6.62 15.21
+34B 28.64 16.80 10.51 17.91
+Deepseek-Coder-Base 6.7B 40.23 22.12 13.04 23.92
+Deepseek-Coder-Instruct 6.7B 44.65 23.86 12.89 25.83
+MoTCoder 6.7B 50.01 29.81 14.36 30.76
+GPT-4 - 34.97 13.75 14.52 18.15
 4.1 Ablation Experiments
-In this section, we perform ablation experiments to examine the effectiveness of our MoTCoder training framework.
-4.2.1 Modular-of-Thought Instruction-Tuning
-We propose two instruction-tuning strategies:
-  1. 1.
-    One-step strategy. The models directly generate the necessary modules and incorporate them into a complete solution.
-  2. 2.
-    Two-step strategy.
-Initially, the instructions outline the essential sub-modules, generating their function headers and docstrings that describe their intended purpose.
-Subsequently, the instruction guides the model to implement these outlined modules and eventually merge them into a cohesive final solution.
-Note that both strategies generate substeps within a single process.
-As indicated in Table 2, the suggested two-step modular-of-thought instruction tuning strategy outperforms the other both for normal and modular inference.
-This aligns with the problem-solving approach employed by adept developers: When faced with a problem, skilled developers design solutions with a modular structure, organizing them into high-level logical sub-modules.
-They subsequently finalize and integrate these components to effectively produce their final solutions.
-4.2.2 Training Dataset
 In this section, we conduct ablation experiments on the dataset, as illustrated in Table 1.
-Initially, we finetuned our model on the modular dataset obtained after modular-of-though instruction evolution, resulting in a substantial improvement (11.0%) compared to the baseline of normal fine-tuning.
-Subsequently, we explore the incorporation of a normal dataset, conducting experiments on a mixture of normal and modular datasets, revealing further performance enhancement of 3.38%.
-This underscores the significance of data diversity: While we encourage the model to decompose tasks into submodules, we also allow the model to employ direct methods to solve problems.
-This provides a larger solution space, leading to significant improvements, especially for easy and medium-level problems (2.46% and 4.54%, respectively).
-The improvement is more modest for difficult problems (0.84%), aligning with the intuition that direct methods are better suited for simpler problems.
-Table 6: CodeContests test and validation results by pass@k (%).
-The rounds of revision after the first round of direct generation is 5 for CodeChain (Le et al., 2023).
-SOTA models includes code-davinci, GPT3.5 and WizardCoder (Luo et al., 2023b), with code-revision method CodeChain (Le et al., 2023) and filtering method CoT(Chen et al., 2023).
-Model Size Revision Validation Test Average
-pass@1 pass@5 pass@1 pass@5 pass@1 pass@5
-code-davinci-002 - - - - 1.00 - 1.00 -
-code-davinci-002 + CodeT - - - - 3.20 - 3.20 -
-WizardCoder 15B - 1.11 3.18 1.98 3.27 1.55 3.23
-WizardCoder + CodeChain 15B 5 2.35 3.29 2.48 3.30 2.42 3.30
-MoTCoder 15B - 2.39 7.69 6.18 12.73 4.29 10.21
-GPT3.5 - - 6.81 16.23 5.82 11.16 6.32 13.70
+We evaluated the performance of instruction tuning on our proposed Module-of-Thought (MoT) dataset versus original dataset, which is a small part of training set of CodeContests and APPS.
+We utilize the DeepSeek-Coder-6.7B model as the base model.
+The results, as delineated in the table below, highlight a pivotal finding: while the MoT tuning strategy maintains comparable performance on introductory tasks, it significantly enhances outcomes in competitive programming scenarios, elevating the pass@1 rate from 11.52% to 14.36%.
+The results align with our intuition that breaking down code into sub-modules is more suitable for complex tasks, whereas simpler tasks are better suited for direct generation.
+The overall enhancement underlines the effectiveness of our MoT method in elevating model proficiency in competitive programming domains.
+Table 3: CodeContests test results by pass@k (%).
+The rounds of revision after the first round of direct generation is 5 for CodeChain Le et al.
+(2023).
+SOTA models include StarCoder Li et al.
+(2023b), WizardCoder Luo et al.
+(2023c), CodeLlama series models Roziere et al.
+(2023), code-davinci, GPT4 OpenAI (2023b), with code-revision method CodeChain Le et al.
+(2023) and filtering method CoTChen et al.
+(2023).
+Model Size pass@1 pass@5
+code-davinci-002 - 1.00 -
+code-davinci-002 + CodeT - 3.20 -
+WizardCoder 15B 1.98 3.27
+WizardCoder + CodeChain 15B 2.48 3.30
+Octocoder 16B 4.95 13.03
+Codellama 7B 0.30 1.11
+13B 2.12 6.26
+34B 5.35 12.02
+Codellama-Python 7B 4.75 10.30
+13B 4.75 12.32
+34B 5.86 14.85
+Codellama-Instruct 7B 2.12 6.26
+13B 5.96 12.02
+34B 6.46 14.24
+Deepseek-Coder-Base 6.7B 6.46 15.25
+Deepseek-Coder-Instruct 6.7B 6.87 8.18
+MoTCoder 6.7B 9.29 16.97
+GPT-4 - 16.36 -
 4.2 Results
-In this section, we present a comprehensive comparison of our approach with existing large coding model baselines on APPS  (Hendrycks et al., 2021a) in Section 4.3.1 and CodeContest (Li et al., 2022a) in Section 4.3.2.
-We employ normal inference prompts for all methods.
+In this section, we present a comprehensive comparison with existing large coding model baselines on APPS  (Hendrycks et al., 2021a) and CodeContest (Li et al., 2022a).
 4.2.1 Results on APPS
@@
 All outcomes are computed using raw predictions without being filtered by the test cases provided in the prompt.
-Our analysis includes a comparison with open-sourced approaches such as CodeT5 (Wang et al., 2021), fine-tuned GPT-Neo (Hendrycks et al., 2021a), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), one-shot StarCoder (Li et al., 2023), and one-shot WizardCoder (Luo et al., 2023b).
-Additionally, we present results from closed-source models, including text-davinci, code-davinci, and GPT3.5.
-As depicted in the results from the APPS shown in Table 4, we notice that the modular-of-thought training approach leads to an enhancement in code generation capabilities compared with previous instruction finetuning models.
+Our analysis includes a comparison with open-sourced approaches such as CodeT5 (Wang et al., 2021), fine-tuned GPT-Neo (Hendrycks et al., 2021a), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), one-shot StarCoder Li et al.
+(2023b), WizardCoder (Luo et al., 2023c), and CodeLlama series (Roziere et al., 2023).
+Additionally, we present results from SOTA closed-source models GPT-4 (OpenAI, 2023b).
+As depicted in the results from the APPS shown in Table 2, we notice that the module-of-thought training approach leads to an enhancement in code generation capabilities compared with previous instruction finetuning models.
 Our MoTCoder exhibits improved performance across all difficulty levels and demonstrates more substantial gains in the interview and competition-level problems, characterized by more intricate solutions.
-To provide specifics, the pass@1 performance of MoTCoder-15B surpasses WizardCoder-15b by 15.49% on interview-level problems and 10.28% on competition-level problems.
-Moreover, our pass@1 on competition-level questions surprisingly outperformed the closed-source model GPT3.5 by an impressive margin of 5.7%.
-This outcome serves as compelling evidence of the efficacy of our approach in addressing intricate problems.
+To provide specifics, the pass@1 performance of MoTCoder-6.7B surprisingly outperformed the closed-source model GPT-4 by an impressive margin of 12.61%.
+This outcome serves as compelling evidence of the efficacy of our approach in addressing competitive programming problems.
 We also conduct a comparative analysis of our approach against previous LLM baselines with code-revision methods as well.
-Our included baselines contain Codex (Chen et al., 2021c), CodeT5 (Wang et al., 2021), code-davinci, StarCoder (Li et al., 2023), and WizardCoder (Luo et al., 2023b) and code-revision methods contain Self-edit (Zhang et al., 2023), CodeRL (Wang et al., 2021; Le et al., 2022), Self-repair (Olausson et al., 2023), and CodeChain (Le et al., 2023).
+Our included baselines contain Codex (Chen et al., 2021c), CodeT5 (Wang et al., 2021), code-davinci, StarCoder (Li et al., 2023b), and WizardCoder (Luo et al., 2023c) and code-revision methods contain Self-edit (Zhang et al., 2023), CodeRL (Wang et al., 2021; Le et al., 2022), Self-repair (Olausson et al., 2023), and CodeChain (Le et al., 2023).
 The results presented in Table 2 illustrate that MoTCoder exhibits notable performance improvements.
-Specifically, MoTCoder achieves a 20.80% pass@1, surpassing the current SOTA, CodeChain + WizardCoder, by 10.3%.
+Specifically, MoTCoder achieves a 30.76% pass@1, surpassing CodeChain + WizardCoder, by 20.26%.
 CodeChain (Le et al., 2023) is an iterative framework for inference that elicits modularized code generation through a chain of self-revisions.
-In contrast to CodeChain, our approach introduces a guided modular-of-thought framework during training, making it more intrinsic.
+In contrast to CodeChain, our approach introduces a guided module-of-thought framework during training, making it more intrinsic.
 This leads to improvements in performance.
-Notably, this achievement is obtained without the need for any additional revision iterations, thereby avoiding any extra associated inference costs.
+Notably, this achievement is obtained without the need for any additional revision iterations, thereby avoiding any extra inference costs.
 4.2.2 Results on CodeContests
-We conduct a comprehensive evaluation of our approach on CodeContests, benchmarking it against current state-of-the-art models, including code-davinci, GPT3.5, and WizardCoder (Luo et al., 2023b).
+We conduct an evaluation of our approach on CodeContests (Li et al., 2022a), benchmarking it against current state-of-the-art models, including code-davinci, GPT3.5, and WizardCoder (Luo et al., 2023c).
 The competitive models utilize the code-revision method CodeChain (Le et al., 2023) and the filtering method CodeT (Chen et al., 2023).
@@
 Specifically, MoTCoder attains an impressive 12.73% pass@5 on the test set, surpassing WizardCoder with CodeChain by 9.43%.
-Remarkably, MoTCoder even outperforms GPT3.5 by 1.57%.
-Importantly, these achievements are realized without the need for additional revision iterations or filtering by test samples, thereby avoiding any additional associated inference costs or the introduction of extra information.
+These achievements are realized without the need for additional revision iterations or filtering by test samples, thereby avoiding any additional associated inference costs or the introduction of extra information.
 5 Limitations
 Through experimentation, it has been observed that although our MoTCoder elevates the model’s modular and logical capabilities, leading to a commendable performance in addressing competitive-level programming challenges, it concurrently sacrifices some of the model’s conversational abilities.
-This finding, along with corroborative evidence from prior research (Roziere et al., 2023), underscores the persistent and crucial nature of addressing the forgetting problem.
+This finding, along with corroborative evidence from prior research Roziere et al.
+(2023), underscores the persistent and crucial nature of addressing the forgetting problem.
 We remain committed to ongoing exploration of this issue in our future work.
 6 Conclusion
-We propose a novel framework, MoTCoder, for tuning instructions based on a modular of thought.
-MoTCoder directs LLMs to decompose their solutions into modular segments, culminating in the synthesis of a comprehensive final solution.
-Our experiments demonstrate that modular-of-thought strategy significantly improves LLM performance, achieving SOTA passing rate on challenging code tasks and showcasing superior modular capabilities.
-It is crucial in large-scale code development environments, since developers commonly leverage code sub-modules to construct code frameworks over crafting each line of code from scratch.
-MoTCoder empowers the creation of entire code structures using only standardized descriptions of functions.
-Adopting this methodology, we unlock the potential to generate extensive and intricate code frameworks.
+This study highlights the limitations of Large Language Models (LLMs) in solving complex programming tasks due to their tendency to generate monolithic code blocks.
+In response, we developed Module-of-Thought Coder (MoTCoder), a novel framework that encourages the breakdown of tasks into manageable sub-tasks and sub-modules.
+Our results demonstrate that MoTCoder’s approach significantly enhances the modularity and accuracy of solutions, as evidenced by considerable improvements in pass@1 rates on both APPS and CodeContests benchmarks.
+We believe the introduction of MoT instruction tuning as a method to cultivate and leverage sub-modules paves the path for a promising direction for future research.
+Appendix A Broader impacts
+Positive societal impacts.
+We propose a method for challenging code programming, which will be helpful for users to solve code problems.
+negative societal impacts.
+This approach may be abused in programming contests that are supposed to be done by humans.
+Appendix B Examples
+For given input problems, we provided submodules and final code solution examples for problems from APPs test dataset.
+B.0.1 APPs Example 1
+Listing 1: Problem 1 from APPs test dataset
+An accordion is a string (yes, in the real world accordions are musical instruments, but lets forget about it for a while) which can be represented as a concatenation of: an opening bracket (ASCII code \$091\$), a colon (ASCII code \$058\$), some (possibly zero) vertical line characters (ASCII code \$124\$), another colon, and a closing bracket (ASCII code \$093\$).
+The length of the accordion is the number of characters in it.
+For example, [::], [:||:] and [:|||:] are accordions having length \$4\$, \$6\$ and \$7\$. (:|:), {:||:}, [:], ]:||:[ are not accordions.
+You are given a string \$s\$.
+You want to transform it into an accordion by removing some (possibly zero) characters from it.
+Note that you may not insert new characters or reorder existing ones.
+Is it possible to obtain an accordion by removing characters from \$s\$, and if so, what is the maximum possible length of the result?
+—–Input—–
+The only line contains one string \$s\$ (\$1 $\le$ |s| $\le$ 500000\$). It consists of lowercase Latin letters and characters [, ], : and |.
+—–Output—–
+If it is not possible to obtain an accordion by removing some characters from \$s\$, print \$-1\$. Otherwise print maximum possible length of the resulting accordion.
+—–Examples—–
+Input
+|[a:b:|]
+Output
+4
+Input
+Output
+-1
+Listing 2: Submodules generated for Problem 1 from APPs test dataset
+def find_colon_indices(s):
+Description: This function finds the indices of the first and last occurrence of ":" in the given string.
+Input:
+s (str): A string containing colons.
+Output:
+ind (int): The index of the first occurrence of ":".
+bind (int): The index of the last occurrence of ":".
+def count_pipes_between_colons(s, ind, bind):
+Description: This function counts the number of "|" characters between the first and last occurrence of ":" in the given string.
+Input:
+s (str): A string containing colons and "|" characters.
+ind (int): The index of the first occurrence of ":".
+bind (int): The index of the last occurrence of ":".
+Output:
+ans (int): The count of "|" characters between the first and last occurrence of ":".
+Listing 3: Code solution generated for Problem 1 from APPs test dataset
+def find_colon_indices(s):
+ind = s.find(":")
+bind = s.rfind(":")
+return ind, bind
+def count_pipes_between_colons(s, ind, bind):
+ans = 0
+if ind!= 0:
+ans += s.count("|", 0, ind)
+if bind!= len(s) - 1:
+ans += s.count("|", bind + 1, len(s))
+return ans
+s = input()
+ind, bind = find_colon_indices(s)
+ans = count_pipes_between_colons(s, ind, bind)
+if ind == -1 or bind == -1 or ind > bind:
+print(-1)
+else:
+print(bind - ind + 1 + ans)
+B.0.2 APPs Example 2
+Listing 4: Problem 2 from APPs test dataset
+Anton has the integer x. He is interested what positive integer, which doesnt exceed x, has the maximum sum of digits.
+Your task is to help Anton and to find the integer that interests him. If there are several such integers, determine the biggest of them.
+—–Input—–
+The first line contains the positive integer x \$(1 $\leq$ x $\leq$ 10^18)\$ - the integer which Anton has.
+—–Output—–
+Print the positive integer which doesnt exceed x and has the maximum sum of digits. If there are several such integers, print the biggest of them. Printed integer must not contain leading zeros.
+—–Examples—–
+Input
+100
+Output
+99
+Input
+48
+Output
+48
+Input
+521
+Output
+499
+Listing 5: Submodules generated for Problem 2 from APPs test dataset
+def find_best_number(num):
+Description: This function finds the best number by decrementing digits from the given number.
+Input:
+num (list): A list of digits representing the number.
+Output:
+best (list): A list of digits representing the best number.
+def convert_to_string(num):
+Description: This function converts a list of digits into a string representation.
+Input:
+num (list): A list of digits.
+Output:
+s (str): The string representation of the digits.
+Listing 6: Code solution generated for Problem 2 from APPs test dataset
+def find_best_number(num):
+best = num[:]
+for i in range(8, -1, -1):
+if num[i] > 0:
+num[i] -= 1
+break
+return num
+def convert_to_string(num):
+s = .join(map(str, num))
+return s
+x = int(input())
+num = [int(i) for i in str(x)]
+best = find_best_number(num)
+s = convert_to_string(best)
+print(s)
+print(s)